OpenStudy (anonymous):

4i(-3-yi)+7=27-12i What is the value of y?

OpenStudy (alekos):

first you need to expand the LHS

OpenStudy (alekos):

4i(-3-yi) = -12i - 4yi^2 = ?

OpenStudy (anonymous):

What's LHS?

OpenStudy (anonymous):

I'm not sure how you did that or how it helps me find the y value but I copied that question verbatim so I can't expand on anything

OpenStudy (alekos):

LHS is the left hand side of the equation which I've just expanded and now I'm asking you to do the next step?

OpenStudy (anonymous):

Is it 4i(3-yi)=-12i+4

OpenStudy (anonymous):

I mean+ 4y

OpenStudy (alekos):

yes it's -12i + 4y so we now have -12i + 4y = 27 - 12i Now solve for y. Add 12i to both sides and what do we get?

OpenStudy (anonymous):

You get 4y=27?

OpenStudy (anonymous):

Then you get y= 6.75?